OpenStudy (anonymous):

Can someone help me make up a quadratic equation that has two real zeros? PLEASE

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

I have no idea how to though

OpenStudy (anonymous):

I know that the formula is \[Ax ^{2} + Bx + C = 0\]

OpenStudy (anonymous):

@3.14159Vikki

OpenStudy (michele_laino):

a possible equation, is: \[\left( {x - a} \right)\left( {x - b} \right) = 0\] where a, and b are the zeroes of that equations, namely they are real numbers

OpenStudy (anonymous):

So I can substitute any number in for x?

OpenStudy (michele_laino):

developing that formula, we get: x^2-(a+b)x+ab=0 so, comparing with your formula, we get: A=1, B=-(a+b), C=a*b

OpenStudy (anonymous):

Thats confusing

OpenStudy (michele_laino):

yes!, since it is a quadratic function of x

OpenStudy (anonymous):

This is what I have to do You are going to write a fairy tale! Your story should include all of the standard elements of fairy tales. Begins “Once upon a time … ” Animals can act like humans. Involves magical items or helpers. Ends “ … happily ever after.” Your story must also fully cover these math concepts. This is an educational fairy tale … the best kind! Create a quadratic function with two real zeros. Solve a system of non-linear equations with a graph. Solve a system of non-linear equations with a table. Compare linear, quadratic, and exponential growth functions using a table and a graph. Submit the graph and table. Identify that exponential functions exceed linear and quadratic functions. You may present your fairy tale as a document, a slideshow presentation, a comic book, a video, or other medium. Any steps or work must be shown.

OpenStudy (michele_laino):

we have: \[\left( {x - a} \right)\left( {x - b} \right) = {x^2} - \left( {a + b} \right)x + ab = 0\]

OpenStudy (anonymous):

How do you get the second part of the equation?

OpenStudy (michele_laino):

I have applied the distributive property of multiplication, over addition

OpenStudy (anonymous):

Oh okay well can you help me with the instructions that I copied and pasted

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

please wait a moment

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

we can consider the subsequent system: \[\left\{ \begin{gathered} {x^2} + {y^2} = 25 \hfill \\ y = 2x - 1 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

@jewelshipp

OpenStudy (anonymous):

I don't know what that means I'm doing algebra 1 online and it doesn't explain things very well

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

in that system we are looking for the intersection between a circumference whose radius is r=5, and whose center is (0,0) namely the origin, and the line whose equation is y=2x-1

OpenStudy (michele_laino):



OpenStudy (anonymous):

I didn't learn anything about circles in this module or lesson

OpenStudy (michele_laino):

do you know the equation of a parabola?

OpenStudy (anonymous):

Ax^2 + Bx + C = 0

OpenStudy (michele_laino):

better is: y=Ax^2+Bx+C

OpenStudy (anonymous):

Okay now i think all its asking is for me to substitute A, B, and C right?

OpenStudy (michele_laino):

yes! Nevertheless in order to apply that procedure, I have to give you three point, at least

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But when you put it in the quadratic equation it has to have two real zeros

OpenStudy (michele_laino):

the number of the real zeroes of a quadratic equation, depends on the value of this quantity: B^2-4*A*C if that quantity is greater than or equal to zero, then we have 2 real zeroes. more precisely if B^2-4*A*C >0, then we have 2 distinct real zeroes, if B^2-4*A*C=0 then we have 2 identical real zeroes whereas if: B^2-4*A*C <0, then we have no real zeroes, namely the 2 zeroes are complex conjugate numbers

OpenStudy (anonymous):

so i do \[B^2-4\times AC <0\]

OpenStudy (anonymous):

oh wait no

OpenStudy (michele_laino):

no we have to request that thie subsequent condition hods: \[\Large {B^2} - 4AC \geqslant 0\]

OpenStudy (michele_laino):

holds*

OpenStudy (anonymous):

okay can i substitute any number or do they have to be certain numbers?

OpenStudy (michele_laino):

you have to have certain numbers, since the quantities: A, B, abd C are parameters, and not variables

OpenStudy (anonymous):

what numbers then?

OpenStudy (michele_laino):

dempends on the exercise that you have to solve. For example if we have: A=1, B=-6, and C=8 we get: \[A{x^2} + Bx + C = {x^2} - 6x + 8\]

OpenStudy (michele_laino):

depends]

OpenStudy (anonymous):

There isn't a certain situation its just a random thing that i have to make

OpenStudy (michele_laino):

please can you give me an example?

OpenStudy (anonymous):

I don't have one

OpenStudy (anonymous):

I just have to make a fairy tail using math

Parth (parthkohli):

Is this what they call a "QH question"?

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

I don't even know what a QH question is

Parth (parthkohli):

Qualified helpers?

OpenStudy (anonymous):

Oh yeah

Parth (parthkohli):

I mean you must have paid to ask this, right?

OpenStudy (anonymous):

yes because i cant find the answer anywhere

OpenStudy (michele_laino):

I think that every question deserves attention @ParthKohli

OpenStudy (anonymous):

All i need is someone to make up a quadratic equation for me

OpenStudy (michele_laino):

I give you an example fro geometry. Is it ok?

OpenStudy (michele_laino):

from*

OpenStudy (anonymous):

I don't think i will get it because i havn't taken geometry

OpenStudy (michele_laino):

I can give you an example from physics, then

OpenStudy (anonymous):

im in 8th grade dude